### v1

# Quantum Shortcut to Adiabaticity for State Preparation in a Finite-Sized Jaynes-Cummings Lattice

Abstract
In noisy quantum systems, achieving high-fidelity state preparation using the adiabatic approach faces a dilemma: either extending the evolution time to reduce diabatic transitions or shortening it to mitigate decoherence effects. Here, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized Jaynes-Cummings lattice by applying a counter-diabatic (CD) driving along given adiabatic trajectories. Leveraging the symmetry of eigenstates in this system, we derive a simplified CD Hamiltonian that only involves local qubit-cavity couplings for a two-site lattice with one polariton excitation. Additionally, we derive the analytical form of the CD Hamiltonian for this lattice with two excitations. Our numerical results demonstrate that this scheme is robust against circuit errors and environmental noise, with characterization achievable through qubit detection. The simplified CD Hamiltonian can be implemented in physical systems with realistic parameters. This approach can lead to a promising pathway to high-fidelity state preparation within a significantly reduced timescale compared to conventional adiabatic methods.
I Introduction
Adiabatic evolution involves a gradual transformation of a quantum system from an initial state to a desired final state by slowly varying a time-dependent Hamiltonian, and has been extensively studied for quantum state preparation in quantum information processing [1, 2]. In noisy quantum systems [3, 4], achieving high fidelity for the prepared states using the adiabatic approach requires a balance in choosing the evolution time: it should be long enough to reduce unwanted diabatic transitions yet short enough to mitigate decoherence from environmental noise. Various approaches have been used to develop quantum shortcut to adiabaticity, where reverse engineering of diabatic transitions is employed to accelerate a slow adiabatic process via nonadiabatic shorcuts [5, 6]. These approaches include counter-diabatic (CD) driving [7, 8, 9, 10, 11], invariant-based approaches [12, 13], the derivative removal of adiabatic gate (DRAG) approach [14], dressed state approaches [15] and etc.. Quantum shortcut to adiabaticity has been proposed and experimentally demonstrated in many quantum systems, such as superconducting qubits, defect qubits, quantum dot arrays, and cold atoms [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. One hurdle that prevents the general application of the quantum shortcut approaches is the complexity of the Hamiltonian required for reverse engineering, which often involves multipartite or nonlocal interactions. Approximate approaches have been explored to overcome this hurdle, including mean-field approximation, variational approach of the effective Hamiltonian, and digitalized approaches [29, 30, 31, 32, 33].
Jaynes-Cummings (JC) lattices [34, 35, 36, 37, 38, 39, 40, 41, 42] have been studied in various experimental systems, including superconducting qubits coupled to cavities, solid-state defects coupled to nanocavities, and trapped ions [43, 44, 45, 46, 47, 48, 49, 50], and can exhibit quantum or dissipative phase transitions between the Mott-insulating and superfluid phases in the thermodynamic limit [51, 52, 40, 41]. The states of finite-sized JC lattices can also carry related features, resulting in novel phenomena such as the photon blockade effect and novel entanglement when properly populated with polariton excitations [53, 54]. A key step in demonstrating these effects is to prepare desired quantum states with a finite number of excitations. In previous works, we employed an adiabatic approach and a quantum optimal control approach for state preparation in this system [55, 56].
In this work, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized JC lattice by applying CD driving during an adiabatic evolution. In our approach, we explore the symmetry of the system’s eigenstates to simplify the CD Hamiltonians for two-site and three-site lattices. We find that for a two-site lattice with one excitation, the CD Hamiltonian can involve only local qubit-cavity couplings, and with two excitations, the CD Hamiltonian on a subspace of the system can have an analytical form involving four-operator couplings. We also study the effects of control errors and environmental noise on this scheme. Our numerical results demonstrate that this scheme is robust against circuit errors and environmental noise and can be characterized by measuring the qubits. This work hence presents an implementable CD Hamiltonian that shortens the evolution time of an adiabatic process in a JC lattice and provides a promising avenue to achieving high-fidelity state preparation within a significantly reduced timescale.
II Jaynes-Cummings Lattice
Consider a finite-sized JC lattice composed of JC models coupled via photon hopping between adjacent sites, as illustrated in Fig. 1. The Hamiltonian of the JC lattice can be written as $H_{\rm JC}=H_{0}+gV_{g}+JV_{J}$, where $H_{0}=\sum_{j}(\omega_{0}a_{j}^{\dagger}a_{j}+\frac{\omega_{z}}{2}\sigma_{jz})$ is the Hamiltonian of the uncoupled qubits and cavity modes, $V_{g}=\sum_{j}(a_{j}^{\dagger}\sigma_{j-}+\sigma_{j+}a_{j})$ describes the onsite qubit-cavity coupling, and $V_{J}=-\sum_{j}(a_{i}^{\dagger}a_{i+1}+a_{i+1}^{\dagger}a_{i})$ describes the photon hopping between neighboring sites. Here, $a_{j}$ ($a_{j}^{\dagger}$) is the annihilation (creation) operator of the cavity modes, $\sigma_{j\pm}$, $\sigma_{jz}$ are the Pauli operators of the qubits, $\omega_{0}$ is the frequency of the cavities, $\omega_{z}$ is the qubit energy splitting, $g$ is the strength of the qubit-cavity coupling, $J$ is the photon hopping rate, and $j\in[1,\,N]$ with $N$ being the number of sites. We have assumed $\hbar=1$ for convenience of discussion. In the rotating frame defined by $H_{0}^{\rm rot}=\omega_{z}\sum_{j}(a_{j}^{\dagger}a_{j}+\frac{1}{2}\sigma_{jz})$, the Hamiltonian of the JC lattice becomes $H_{\rm r}=\sum_{j}\Delta a_{j}^{\dagger}a_{j}+gV_{g}+JV_{J}$, where $\Delta=\omega_{0}-\omega_{z}$ is the detuning between the cavity modes and the qubits.
Figure 1: Schematic of a two-site JC lattice with (a) one polariton excitation and (b) two polariton excitations. Each site is composed of a qubit (two-level system) coupled to a cavity mode with coupling strength $g$, and the cavities are connected by photon hopping with rate $J$. In (a), the CD driving includes only local qubit-cavity couplings denoted by $g^{\prime}$; in (b), the CD driving is shown as $H_{1}^{\prime}$.
We denote the basis states of a single JC model as $|n,s\rangle$, with $n\geq 0$ being the photon number of the cavity mode and $s=g,\,e$ being the ground and excited states of the qubit, respectively. The eigenstates of a single JC model include the ground state $|0,g\rangle$ and the doublets $|n,\pm\rangle$ ($n\geq 1$), which are superpositions of the basis states $|n,g\rangle$ and $|n-1,e\rangle$ and correspond to eigenstates with $n$ polariton excitations. In the thermodynamic limit with $N\rightarrow\infty$, a JC lattice can exhibit various many-body effects such as the quantum and dissipative phase transitions between the Mott insulating and the superfluid phases [34, 35, 36]. It can be shown that the ground states of a finite-size JC lattice exhibit similar features to these phases [40, 55, 56].
A key step in studying novel phenomena in a JC lattice is the preparation of desired quantum states for a finite number of polariton excitations. One approach to generate such states is adiabatic evolution [55], where the system is first prepared in the ground state of an initial Hamiltonian $H_{\rm r}(t_{0})$ at time $t_{0}$, followed by adiabatic tuning of the system parameters to reach the target Hamiltonian $H_{\rm r}(T)$ at the final time $T$. The ground state of the target Hamiltonian is designed to be the desired state, which is often hard to prepare directly by gate operations. In the end of the adiabatic evolution, the system will reach the desired state. However, the adiabaticity of the evolution and hence the fidelity of the prepared states can be impaired by unwanted diabatic transitions to excited states. Below, we will employ a quantum shortcut to adiabaticity approach to study the robust generation of quantum states in a JC lattice within a significantly reduced timeframe.
III Counter-Diabatic Driving
CD driving can be applied to the Hamiltonian $H_{\rm r}(t)$ to cancel diabatic transitions and generate quantum shortcut to adiabaticity during an adiabatic evolution. The CD Hamiltonian has the general form [7, 8, 5]:
$H_{1}(t)=i\sum_{m\neq n}\frac{|m(t)\rangle\langle m(t)|\partial_{t}H_{\rm r}|n(t)\rangle\langle n(t)|}{E_{n}(t)-E_{m}(t)},$ (1)
where $|m\rangle$, $|n\rangle$ are instantaneous eigenstates of $H_{\rm r}(t)$ at time $t$ and $E_{m}$, $E_{n}$ are the corresponding eigenenergies. The total Hamiltonian of this system then becomes $H_{\rm tot}=H_{\rm r}(t)+H_{1}(t)$. The exact CD Hamiltonian in (1) often includes nonlocal or multipartite interactions due to the complexity of the quantum system of interest. Hence, it is challenging to implement this Hamiltonian in systems with practical parameters.
In this section, we derive and simplify the CD Hamiltonian for two-site and three-site JC lattices evolving through given adiabatic trajectories. Our results show that the simplified CD Hamiltonian can involve only local qubit-cavity coupling and can be implementable using current technology.
III.1 Two sites one excitation
Eigenstates. We first study a two-site JC lattice ($N=2$) with only one polariton excitation. The allowable Hilbert space for this system includes four basis states: $|1,g\rangle_{1}|0,g\rangle_{2}$, $|0,e\rangle_{1}|0,g\rangle_{2}$, $|0,g\rangle_{1}|1,g\rangle_{2}$, and $|0,g\rangle_{1}|0,e\rangle_{2}$, where the subscripts refer to the sites 1 and 2 in the lattice. Written in terms of this basis set, the Hamiltonian has the form:
$H_{\rm r}=\begin{pmatrix}\Delta&g&-J&0\\ g&0&0&0\\ -J&0&\Delta&g\\ 0&0&g&0\end{pmatrix}$ (2)
for given values of $g,\,J,$ and $\Delta$. The eigenvalues $E_{n}$ ($n\in[1,4]$) of this system are
$E_{1,2}=\frac{1}{2}(\Delta_{J}^{-}\mp\chi_{1}^{-}),\quad E_{3,4}=\frac{1}{2}(\Delta_{J}^{+}\mp\chi_{1}^{+})$ (3)
with $\Delta_{J}^{\pm}=\Delta\pm J$ and $\chi_{1}^{\pm}=\sqrt{(\Delta_{J}^{\pm})^{2}+4g^{2}}$. The corresponding eigenvectors are
$v_{1}$ $=\frac{1}{2}\begin{pmatrix}-\sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ -\sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\end{pmatrix},\quad v_{2}=\frac{1}{2}\begin{pmatrix}\sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\end{pmatrix},$ (4)
$v_{3}$ $=\frac{1}{2}\begin{pmatrix}\sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ -\sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ -\sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ \sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\end{pmatrix},\quad v_{4}=\frac{1}{2}\begin{pmatrix}-\sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ -\sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ \sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ \sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\end{pmatrix}.$
The eigenstate $v_{1}$ is the ground state with energy $E_{1}$. From (4), we observe that the states $v_{1}$ and $v_{2}$ are symmetric with respect to an exchange of the states on the two sites, while the states $v_{3}$ and $v_{4}$ are antisymmetric and acquire a factor of $-1$ when the states on the two sites are swapped.
Figure 2: Eigenenergies $E_{n}$ of a two-site JC lattice with one excitation. (a) $E_{n}$ vs hopping rate $J$ for $g=1$ and $\Delta=1$. (b) $E_{n}$ vs qubit-cavity coupling $g$ for $J=2$ and $\Delta=1$. All parameters are given in dimensionless units.
In Fig. 2(a), we plot the energy spectrum of the system as a function of the photon hopping rate $J$ for $g=1$ and $\Delta=1$. At $J=0$, $\Delta_{J}^{\pm}=\Delta$, resulting in $E_{1}=E_{3}$ ($E_{2}=E_{4}$), which indicates that the states $v_{1}$ and $v_{3}$ ($v_{2}$ and $v_{4}$) are degenerate states for arbitrary values of the qubit-cavity coupling and detuning. In Fig. 2(b), we plot the energy spectrum vs the qubit-cavity coupling $g$ for $J=2$ and $\Delta=1$. At $g=0$, $E_{2}=E_{3}=0$ when $\Delta<J$.
We want to emphasize that the symmetry of these eigenstates is crucial for our development of a simplified CD Hamiltonian that exclusively involves local qubit-cavity couplings, making it feasible for implementation in practical systems. This symmetry also facilitates the preparation of the ground state $v_{1}$ at $J=0$ via resonant pumping, without inducing excitation to the degenerate state $v_{3}$. Furthermore, during an adiabatic evolution, the variation of the Hamiltonian does not transition the system from a symmetric to an antisymmetric state due to the inherent symmetry of the Hamiltonian.
Ramping of hopping rate $J$. Assuming that the hopping rate $J(t)$ is continuously tuned from $J(0)=0$ at time $t=0$ to $J(T)=J_{f}$ at the final time $T$ with $J_{f}$ being the target hopping rate, while keeping the qubit-cavity coupling and the detuning unchanged during the adiabatic process. At $t=0$, the system is prepared in the ground state $v_{1}$ for $J=0$ by resonantly pumping the system from the state $|0,g\rangle_{1}|0,g\rangle_{2}$. The variation of $J(t)$ induces off-diagonal matrix elements between the instantaneous eigenstates, resulting in undesired transitions to the excited states. The nonzero matrix elements are: $|\langle v_{1}|V_{J}|v_{2}\rangle|=g/\chi_{1}^{-}$, $|\langle v_{3}|V_{J}|v_{4}\rangle|=g/\chi_{1}^{+}$, along with their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$, and between the states $v_{3}$ and $v_{4}$, respectively. We then derive the CD Hamiltonian using (1):
$H_{1}(t)$ $=$ $i\frac{g}{(\chi_{n}^{-})^{2}}\frac{dJ}{dt}(|v_{1}\rangle\langle v_{2}|-|v_{2}\rangle\langle v_{1}|)$ (5)
$-i\frac{g}{(\chi_{n}^{+})^{2}}\frac{dJ}{dt}(|v_{3}\rangle\langle v_{4}|-|v_{4}\rangle\langle v_{3}|).$
Hence, when the system is in the ground state $|v_{1}\rangle$, the only allowable transition is to the symmetric eigenstate $|v_{2}\rangle$. As discussed above, this is due to the symmetry of the eigenstates and the Hamiltonian. A variation of the hopping rate in the Hamiltonian preserves the symmetry of the state, with the only possible diabatic transition being the excitation to the state $|v_{2}\rangle$. Consequently, the state at an arbitrary time $t$ during the evolution remains symmetric and can only be a superposition of the instantaneous $|v_{1}\rangle$ and $|v_{2}\rangle$. Below, we will simplify the CD Hamiltonian using this property.
We first convert the CD Hamiltonian (5) to physical operators of the qubits and cavities in order to discuss its implementation. In the Hilbert space for one polariton excitation, we have $a_{1}^{\dagger}\sigma_{1-}=|1,g\rangle_{1}\langle 0,e|\otimes|0,g\rangle_{2}\langle 0,g|$, $a_{1}^{\dagger}a_{2}=|1,g\rangle_{1}\langle 0,g|\otimes|0,g\rangle_{2}\langle 1,g|$, and so forth. Using (4), we derive:
$|v_{1}\rangle\langle v_{2}|-|v_{2}\rangle\langle v_{1}|=$ $\frac{1}{2}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right),$ (6a)
$|v_{3}\rangle\langle v_{4}|-|v_{4}\rangle\langle v_{3}|=$ $\frac{1}{2}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right),$ (6b)
with $A_{\pm}=a_{1}\pm a_{2}$ and $S_{\pm}=\sigma_{1-}\pm\sigma_{2-}$. The CD Hamiltonian can then be written as
$H_{1}(t)$ $=$ $i\frac{g}{2(\chi_{n}^{-})^{2}}\frac{dJ}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (7)
$-i\frac{g}{2(\chi_{n}^{+})^{2}}\frac{dJ}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
This Hamiltonian comprises local couplings between qubits and cavities on the same site, such as $\sigma_{1+}a_{1}$ and $\sigma_{2+}a_{2}$, as well as nonlocal couplings between qubits and neighboring cavities, such as $\sigma_{1+}a_{2}$ and $\sigma_{2+}a_{1}$.
As seen from (5), when the initial state is the ground state $v_{1}$, the CD Hamiltonian only eliminates excitation from $v_{1}$ to $v_{2}$. Thus, the second term in $H_{1}(t)$ does not actively contribute to the elimination of diabatic transitions, and its amplitude can be adjusted at will without affecting the adiabatic evolution. We then change the coefficient of the second term in (7) from $-ig/2(\chi_{n}^{+})^{2}$ to $ig/2(\chi_{n}^{-})^{2}$, and obtain the simplified CD Hamiltonian
$H_{1}^{\prime}(t)=ig^{\prime}\left(a_{1}\sigma_{1}^{+}-a_{1}^{\dagger}\sigma_{1}^{-}+a_{2}\sigma_{2}^{+}-a_{2}^{\dagger}\sigma_{2}^{-}\right).$ (8)
The Hamiltonian $H_{1}^{\prime}$ includes only local qubit-cavity couplings with a strength given by $g^{\prime}=\frac{g}{(\chi_{n}^{-})^{2}}\frac{dJ}{dt}$, which can be interpreted as a Jaynes-Cummings coupling with a $e^{i\pi/2}$ phase. This Hamiltonian can be engineered by adjusting the local qubit-cavity couplings.
For demonstration, we simulate the time evolution of this system under the adiabatic Hamiltonian $H_{\rm r}$ with either the CD Hamiltonian $H_{1}$ or $H_{1}^{\prime}$ applied. We investigate both linear ramping of the hopping rate, with $J(t)=J_{f}(t/T)$, and quadratic ramping, with $J(t)=J_{f}(t/T)^{2}$. For linear ramping, $dJ/dt=J_{f}/T$; and for quadratic ramping, $dJ/dt=2J_{f}t/T^{2}$. We choose the target hopping rate $J_{f}=2$, the qubit-cavity coupling $g\equiv 1$, the detuning $\Delta\equiv 1$, and the total time $T=0.5\pi$. We define the fidelity of the quantum state at time $t$ as $F(t)=|\langle v_{1}(t)|v(t)\rangle|^{2}$, where $|v(t)\rangle$ is the state at time $t$ and $|v_{1}(t)\rangle$ is the instantaneous ground state of the adiabatic Hamiltonian $H_{\rm r}$ at time $t$. In Fig. 3(a), we plot the infidelity $1-F(t)$ vs $t/T$. Without the CD Hamiltonian, the infidelity increases significantly during the evolution, as $T=0.5\pi$ is a very short time and diabatic transitions can be serious. In contrast, when the CD Hamiltonian is applied, the infidelity remains below $10^{-10}$ throughout the evolution, preserving the system in the ground state. Both CD Hamiltonians $H_{1}$ and $H_{1}^{\prime}$ yield identical results, confirming our analysis. In Fig. 3(b), we plot the infidelity $1-F(T)$ of the final state vs the total evolution time $T$. For short evolution times, the infidelity of the adiabatic process is quite high. However, as $T$ increases, the infidelity decreases to below $10^{-3}$. These numerical results demonstrate that the CD approach can ensure high fidelity for short evolution times, which is crucial for quantum devices in the NISQ regime.
Figure 3: (Color online) (a) Infidelity $1-F(t)$ vs $t/T$ for a total evolution time of $T=0.5\pi$. (b) Infidelity $1-F(T)$ vs $T$. All solid lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; and all circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied. Blue (red): linear (quadratic) ramping of the hopping rate $J$ for $g\equiv 1$, $\Delta\equiv 1$, $J(0)=0$ and $J_{f}=2$; and green: linear ramping of the qubit-cavity coupling $g$ for $J\equiv 2$, $\Delta\equiv 1$, $g(0)=0$ and $g_{f}=1$.
Ramping of qubit-cavity coupling $g$. Now consider an adiabatic trajectory where the qubit-cavity coupling is ramped linearly with $g(t)=g_{f}t/T$ and $g_{f}$ being the target coupling strength. The nonzero transition matrix elements caused by the variation of $g$ are $|\langle v_{1}|V_{g}|v_{2}\rangle|=\Delta_{J}^{-}/\chi_{1}^{-}$, $|\langle v_{3}|V_{g}|v_{4}\rangle|=\Delta_{J}^{+}/\chi_{1}^{+}$, and their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$, as well as between the states $v_{3}$ and $v_{4}$, due to the symmetry of the eigenstates. The CD Hamiltonian can be derived as:
$H_{1}(t)$ $=$ $i\frac{\Delta_{J}^{-}}{2(\chi_{n}^{-})^{2}}\frac{dg}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (9)
$-i\frac{\Delta_{J}^{+}}{2(\chi_{n}^{+})^{2}}\frac{dg}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
Similar to (7), this Hamiltonian includes couplings between qubits and their neighboring cavities. As discussed above, when $|v_{1}\rangle$ is the initial state, only the state $|v_{2}\rangle$ could be excited by diabatic transitions. Therefore, we can vary the coefficient of the second term in (9). By changing this coefficient from $-i\Delta_{J}^{+}/2(\chi_{n}^{+})^{2}$ to $i\Delta_{J}^{-}/2(\chi_{n}^{-})^{2}$, we obtain a simplified CD Hamiltonian:
$H_{1}^{\prime}(t)=ig^{\prime}\left(a_{1}\sigma_{1}^{+}-a_{1}^{\dagger}\sigma_{1}^{-}+a_{2}\sigma_{2}^{+}-a_{2}^{\dagger}\sigma_{2}^{-}\right),$ (10)
which only involves local qubit-cavity couplings with strength $g^{\prime}=\frac{\Delta_{J}^{-}}{(\chi_{1}^{-})^{2}}\frac{dg}{dt}$. Consequently, implementing both the adiabatic Hamiltonian and the CD Hamiltonian in this trajectory only requires tuning the local qubit-cavity couplings. This feature can significantly simplify the experimental setup.
We perform a numerical simulation of this process, choosing a hopping rate of $J\equiv 2$, a detuning of $\Delta\equiv 1$, and a target coupling strength of $g_{f}=1$. At $t=0$, with $g(0)=0$, the initial state is $(|1,g\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,g\rangle_{2})/\sqrt{2}$, where the excitation is stored in the cavity modes. The infidelity $1-F(t)$ vs $t/T$ for $T=0.5\pi$ is plotted in Fig.3(a), and the infidelity at the final time $1-F(T)$ vs the total evolution time $T$ is plotted in Fig.3(b). The results exhibit similar behavior to the ramping of the hopping rate $J$.
Self-protected trajectory. The magnitude of the CD Hamiltonian in (10) is proportional to $\Delta_{J}^{-}=\Delta-J$. In the special case of $\Delta=J$, $\Delta_{J}^{-}=0$, leading to the interesting result of $H_{1}^{\prime}=0$. This indicates no CD Hamiltonian is needed to eliminate the diabatic transitions. It can be shown that the matrix element $|\langle v_{1}|V_{g}|v_{2}\rangle|\equiv 0$ under this condition, thus, the only allowable diabatic transition from the ground state to the excited states disappears. The Hamiltonian $H_{\rm r}$ is hence self-protected from diabatic transitions, meaning there is no diabatic transition to the excited states caused by ramping the system parameters. The ground state remains in the state $|v_{1}\rangle=\frac{1}{\sqrt{2}}\left(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2}\right)$ throughout the entire evolution, with the excitation occupying the lower polariton state $|1,-\rangle$ in one of the two lattice sites. This is confirmed by our numerical simulation of this trajectory.
III.2 Three sites one excitation
[…]
$H_{\rm r}=\begin{pmatrix}\Delta&g&-J&0&0&0\\ g&0&0&0&0&0\\ -J&0&\Delta&g&-J&0\\ 0&0&g&0&0&0\\ 0&0&-J&0&\Delta&g\\ 0&0&0&0&g&0\end{pmatrix}.$ (11)
[…]
$H_{1}$ $=$ $ig_{n}\left[\left(a_{1}^{\dagger}+a_{3}^{\dagger}\right)\sigma_{2-}+a_{2}^{\dagger}\left(\sigma_{1-}+\sigma_{3-}\right)+h.c.\right]$ (12)
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}+h.c.\right)$
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}+h.c.\right)$
which includes nonlocal couplings with strength $g_{n}$ or $g_{l}$ between different sites, such as the coupling between the cavity on sites $1$ and $3$ and the qubit on site $2$, as well as local couplings with strengths $g_{l}$ for sites $1$ and $3$, and strength $2g_{l}$ for site $2$.
We employ the same approach as in Sec. III.1 to simplify the Hamiltonian (12). As discussed earlier, when the initial state is $|v_{1}\rangle$, the ramping of the hopping terms induces transitions only to the state $|v_{4}\rangle$. Therefore, the coefficient $\alpha_{2}$ in $H_{1}$ can be varied with flexibility without affecting the evolution of the quantum state. By adjusting $\alpha_{2}$, we derive a simplified CD Hamiltonian in which the couplings between the qubit (cavity) on site $2$ and the cavities (qubits) on sites $1$ and $3$ are no longer required. The simplified CD Hamiltonian takes the form:
$H_{1}^{\prime}$ $=$ $ig^{\prime}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}+h.c.\right)$ (13)
$+ig^{\prime}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}+h.c.\right),$
which includes nonlocal terms between sites $1$ and $3$ with coupling strength $g^{\prime}$. This Hamiltonian can be constructed for a small lattice of three sites by connecting the qubits and cavities on sites $1$ and $3$. However, this result shows that we cannot always simplify the CD Hamiltonian to only contain local couplings
III.3 Two sites two excitations
Eigenstates. For a two-site JC lattice with two polariton excitations, the Hilbert space includes eight basis states: $|2,g\rangle_{1}|0,g\rangle_{2}$, $|1,e\rangle_{1}|0,g\rangle_{2}$, $|0,g\rangle_{1}|2,g\rangle_{2}$, $|0,g\rangle_{1}|1,e\rangle_{2}$, $|1,g\rangle_{1}|1,g\rangle_{2}$, $|0,e\rangle_{1}|1,g\rangle_{2}$, $|1,g\rangle_{1}|0,e\rangle_{2}$, and $|0,e\rangle_{1}|0,e\rangle_{2}$. In the first four states, both excitations occupy one of the two sites, while in the last four states, one excitation is located on each site. Using this basis set, the Hamiltonian $H_{\rm r}$ can be expressed as:
$H_{\rm r}=\begin{pmatrix}2\Delta&\sqrt{2}g&0&0&-\sqrt{2}J&0&0&0\\ \sqrt{2}g&\Delta&0&0&0&-J&0&0\\ 0&0&2\Delta&\sqrt{2}g&-\sqrt{2}J&0&0&0\\ 0&0&\sqrt{2}g&\Delta&0&0&-J&0\\ -\sqrt{2}J&0&-\sqrt{2}J&0&2\Delta&g&g&0\\ 0&-J&0&0&g&\Delta&0&g\\ 0&0&0&-J&g&0&\Delta&g\\ 0&0&0&0&0&g&g&0\end{pmatrix}.$ (14)
For detuning $\Delta=0$, the eigenvalues of this Hamiltonian can be derived analytically as:
$E_{1,2}$ $=0,$ (15a)
$E_{3,4}$ $=\mp\sqrt{2g^{2}+J^{2}},$ (15b)
$E_{5,6}$ $=\mp\frac{\sqrt{6g^{2}+5J^{2}-\sqrt{4g^{4}+60g^{2}J^{2}+9J^{4}}}}{\sqrt{2}},$ (15c)
$E_{7,8}$ $=\mp\frac{\sqrt{6g^{2}+5J^{2}+\sqrt{4g^{4}+60g^{2}J^{2}+9J^{4}}}}{\sqrt{2}}.$ (15d)
In Fig. 4(a), we plot the eigenenergies vs the hopping rate $J$ at $g=1$ and $\Delta=0$. We denote the eigenstates as $|v_{i}\rangle$, with $i\in[1,8]$, corresponding to the eigenvalues given in (III.3). Here, $|v_{7}\rangle$ is the ground state, while $|v_{1}\rangle$ and $|v_{2}\rangle$ are degenerate states with $E_{1,2}=0$.
Figure 4: (a) Eigenenergies $E_{n}$ vs hopping rate $J$. (b) Infidelity $1-F(t)$ vs $t/T$ for total evolution time $T=0.5\pi$. All solid lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; and all circles: adiabatic evolution with the CD Hamiltonian $H_{1}$ or $H_{1}^{\prime}$ applied. Blue: the initial state is the ground state $|v_{7}\rangle$ at $J=0$, and the CD Hamiltonian is $H_{1}$ for the full Hilbert space; red (green): the initial state is the ground state $|v_{7}\rangle$ (the first excited state $|v_{3}\rangle$) at $J=0$, and the CD Hamiltonian is $H_{1}^{\prime}$ for the subspace of $\{|v_{2}\rangle$, $|v_{3}\rangle$, and $|v_{4}\rangle\}$. Here, $g\equiv 1$, $\Delta\equiv 0$, $J(0)=0$, and $J_{f}=1$. The blue solid line is obscured completely by the green solid line.
Ramping of hopping rate $J$. Assuming that the hopping rate is linearly ramped with $J=J_{f}t/T$, the target hopping rate $J_{f}=1$, $g\equiv 1$, and $\Delta\equiv 0$. The CD Hamiltonian for this process can be obtained numerically using Eq.(1). We simulate the dynamics of this system with the initial state being the ground state $|v_{7}\rangle$ for $J=0$ and with a total time $T=0.5\pi$. In Fig. 4(b), we plot the infidelity $1-F(t)$ of the quantum state at time $t$ as a function of $t/T$. The infidelity at the final time $T$ is above $0.4$ for the adiabatic evolution under the Hamiltonian $H_{\rm r}$, whereas it remains negligible when the CD Hamiltonian is applied.
By examining the eigenstates of $H_{\rm r}$, we find that the eight eigenstates can be divided into two subsets. The first subset contains three states: ${|v_{2}\rangle,\,|v_{3}\rangle,\,|v_{4}\rangle}$, while the second subset contains the remaining five states, including the ground state $|v_{7}\rangle$. When varying the hopping rate $J$, the only nonzero transition matrix elements are between states within the same subset; there are no diabatic transitions between states from different subsets. When the initial state is from the first subset, the relevant diabatic transitions occur only within that subset. Thus, the CD Hamiltonian can be simplified by retaining only the counter-diabatic terms for the first subset in $H_{1}$, leading to the construction of a new CD Hamiltonian with these terms. Using this approach, we obtain a simplified CD Hamiltonian:
$H_{1}^{\prime}(t)=-i\frac{g}{2\chi_{2}^{2}}\frac{dJ}{dt}\left(A_{2}^{\dagger}C_{2}-C_{2}^{\dagger}A_{2}\right),$ (16)
where $A_{2}=a_{1}^{2}-a_{2}^{2}$ and $C_{2}=a_{2}\sigma_{1-}-a_{1}\sigma_{2-}$. This Hamiltonian consists of four-operator terms, which can be formed through effective second-order couplings in the perturbative regime. While the engineering of such couplings has been previously studied, it’s important to note that the effective coupling strength may be significantly smaller compared to that of the first-order couplings $g$ and $J$.
To demonstrate the effectiveness of Eq. (16), we conducted a numerical simulation of the adiabatic process using different initial states. In Fig. 4(b), when the initial state is $|v_{3}\rangle$ from the first subset, which corresponds to the first excited state as shown in Fig. 4(a), the infidelity remains below $10^{-10}$. This indicates that the diabatic transitions are completely cancelled by the CD Hamiltonian, preserving the system in the instantaneous eigenstate $|v_{3}\rangle$ throughout the evolution. However, when the initial state is $|v_{7}\rangle$ from the second subset, corresponding to the ground state of the system, the infidelities for the evolution with and without the CD terms are identical, suggesting that the CD Hamiltonian $H_{1}^{\prime}$ cannot cancel the diabatic transitions within the second subset. These numerical findings are consistent with our analysis of the eigenstates and the CD Hamiltonian.
IV Error, Noise, and Measurement
In this section, we discuss the robustness of the CD driving against control errors and environmental noise, as well as the characterization of the CD approach through qubit measurements.
Control errors. Adjusting the system parameters can introduce classical control errors, potentially impacting the performance of the CD approach. To investigate the effects of these errors, we conducted numerical simulations of the adiabatic process by introducing random fluctuations to the time-dependent coupling constants. Consider the linear ramping of the hopping rate $J(t)$ discussed in Sec. III.1, where the CD Hamiltonian only consists of time-dependent local qubit-cavity couplings. We introduced a fluctuation $\delta J(t)$ to the hopping rate, modifying it as $J(t)\rightarrow J(t)+\delta J(t)$, and a fluctuation $\delta g(t)$ to the qubit-cavity coupling [including the CD terms in (10)] with $g(t)\rightarrow g(t)+\delta g(t)$. Here, $\delta J(t)$ and $\delta g(t)$ are Gaussian random numbers with a standard deviation $\alpha J_{f}$ and $\alpha g$, respectively, with $\alpha$ being the ratio of the fluctuation magnitude to the coupling strength. We choose $\alpha$ within the range of $[0.01,0.15]$, i.e., the fluctuations are within $1\%$ $15\%$ of the coupling strengths. In Fig. 5(a), the fidelity $F(t)$ is plotted vs $t/T$ for a single sample of errors with $\alpha=0.05$. The fidelity exhibits small fluctuations over time due to the presence of errors. In Fig. 5(b), the fidelity $F(T)$ at time $T$, averaged over $100$ samples of errors, is plotted vs $\alpha$. The numerical result shows only a mild decrease in fidelity with the increase of $\alpha$, even for $\alpha\sim 15\%$, demonstrating the robustness of this process against control errors.
Figure 5: (a) Fidelity $F(t)$ vs $t/T$ for one sample of errors with $\alpha=0.05$. (b) Fidelity $F(T)$ averaged over 100 samples vs $\alpha$. In (a) and (b), all blue circles: adiabatic evolution under the Hamiltonian $H_{\rm r}$; all red circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied. The total evolution time $T=0.5\pi$. (c) Infidelity $1-F(T)$ vs $\gamma$ for $\kappa=5\times 10^{-5}$. (d) Infidelity $1-F(T)$ vs $\kappa$ for $\gamma=\frac{5}{\pi}\times 10^{-5}$. In (c) and (d), all solid lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; and all circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied. Blue: $T=0.5\pi$; red: $T=3\pi$; and green: $T=5.5\pi$. All other parameters are the same as in Fig. 3.
Decoherence. Environmental noise induces decoherence in the qubits and cavity modes, potentially impacting the fidelity of the prepared quantum states. To analyze the effects of decoherence in this system, we utilize a master equation approach. The master equation takes the form
$\frac{d\rho}{dt}=-i\left[H_{\rm tot},\rho\right]+\sum_{j}\left(\gamma_{j}\mathcal{L}_{qj}+\kappa_{j}\mathcal{L}_{aj}\right)\rho,$ (17)
where $\mathcal{L}_{qj}=\frac{1}{2}(2\sigma_{j-}\rho\sigma_{j+}-\rho\sigma_{j+}\sigma_{j-}-\sigma_{j+}\sigma_{j-}\rho)$ is the Liouvillian operator for the qubit on site $j$ with damping rate $\gamma_{j}$, $\mathcal{L}_{aj}=\frac{1}{2}(2a_{j}\rho a_{j}^{\dagger}-\rho a_{j}^{\dagger}a_{j}-a_{j}^{\dagger}a_{j}\rho)$ is the Liouvillian operator for the cavity mode on site $j$ with damping rate $\kappa_{j}$, and $H_{\rm tot}$ is the total Hamiltonian of the system. Also, we choose $\gamma_{1,2}=\gamma$ and $\kappa_{1,2}=\kappa$ for simplicity of discussion. Assume that the dimensionless coupling strength $g=1$ in our discussion corresponds to $g=2\pi\times 100$ MHz in superconducting systems [55, 56]. The couplings in this system can be controlled by various methods, as demonstrated in recent experiments [57, 58, 59, 60, 61, 62, 63]. The measured decoherence time of $\sim 100{\rm\mu s}$ then corresponds to $\gamma=\frac{5}{\pi}\times 10^{-5}$ in dimensionless units. When the cavity frequency is $\omega_{0}=2\pi\times 5$ GHz and the quality factor is $Q=10^{6}$, the cavity decay rate is $\kappa=5\times 10^{-5}$ in dimensionless units.
[…]
Figure 6: Operator average $\langle\sigma_{1x}\sigma_{2x}\rangle$ vs $t/T$ for $T=0.5\pi$. (a) One excitation with $g\equiv 1$, $\Delta\equiv 1$, and $J_{f}=2$ as in Fig. 3(a). (b) Two excitations with $g\equiv 1$, $\Delta\equiv 0$, and $J_{f}=1$ as in Fig. 4(a). Solid blue lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; blue circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied; and solid red lines: instantaneous ground state average of Hamiltonian $H_{\rm r}$.
Measurement. The performance of the CD approach can be characterized by conducting measurement on the qubit operator $\sigma_{1x}\sigma_{2x}$. In the case of one excitation in a two-site lattice, the ground state at the initial parameters $g=1$, $J=0$, and $\Delta=1$ is $\frac{1}{2}(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2})$, where $|1,-\rangle$ is the lower-polariton state with one excitation, and the excitation occupies the sites in the form of a quantum superposition with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0.7236$. When the hopping rate becomes $J_{f}=2$ at time $T$, the operator average of the ground state decreases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.2764$, as shown in Fig. 6(a). It is noticeable that the operator average for the adiabatic evolution under $H_{\rm r}$ can significantly deviate from that of the ground state, while the operator average with the CD term applied remains the same as that of the ground state. Similarly, in the two-excitation case, the ground state at $g=1$, $J=0$, and $\Delta=0$ is $v_{7}=|1,-\rangle_{1}|1,-\rangle_{2}$, where each site is occupied by one excitation in the lower-polariton state with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0$. When the hopping rate becomes $J_{f}=1$, the operator average increases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.3659$. During the evolution, the operator average for the adiabatic process without the CD driving shows a large discrepancy with the operator average in the CD approach, as shown in Fig. 6(b). Hence, the effectiveness of the CD approach can be verified by measuring the operator average.
V Conclusions
To conclude, we have studied a quantum shortcut to adiabaticity using the counter-diabatic driving approach in a finite-sized JC lattice, leveraging the symmetry inherent in its eigenstates. Our analysis reveals that for a two-site lattice with one excitation, the CD Hamiltonian can be simplified to only involve local qubit-cavity couplings, realizable in practical devices. For more complex scenarios, such as a two-site lattice with two excitations or a three-site lattice with one excitation, the CD Hamiltonians can still be simplified, albeit involving nonlocal terms. Our numerical simulations demonstrate the robustness of the scheme against both control errors and environmental noise and the scheme can be characterized through measurements on the qubits. Overall, our findings offer a promising avenue for achieving high-fidelity quantum state preparation and hold potential for applications across various systems exhibiting similar symmetries.
Acknowledgements
This work is supported by the NSF Award No. 2037987 and the UC-MRPI Program (Grant ID M23PL5936). A.G. is also supported by the Sandbox AQ Fellowship.