### v2

# Quantum Shortcut to Adiabaticity for State Preparation in a Finite-Sized Jaynes-Cummings Lattice

Abstract
In noisy quantum systems, achieving high-fidelity state preparation using the adiabatic approach faces a dilemma: either extending the evolution time to reduce diabatic transitions or shortening it to mitigate decoherence effects. Here, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized Jaynes-Cummings lattice by applying counter-diabatic (CD) driving along given adiabatic trajectories. Leveraging the symmetry of eigenstates in our system, we convert the CD driving to an implementable Hamiltonian that only involves local qubit-cavity couplings for a two-site lattice with one polariton excitation. Additionally, we derive a partial analytical form of the CD driving for the lattice with two excitations. Our numerical results demonstrate that circuit errors and environmental noise have negligible effects on our scheme under practical parameters. We also show that our scheme can be characterized through the detection of qubit operators. This approach can lead to a promising pathway to high-fidelity state preparation in a significantly reduced timescale when compared to conventional adiabatic methods.
I Introduction
Adiabatic evolution involves a gradual transformation of a quantum system from an initial state to a desired final state by slowly varying a time-dependent Hamiltonian, and has been extensively studied for quantum state preparation in quantum information processing [1, 2]. In noisy quantum systems [3, 4], achieving high fidelity for the prepared states using the adiabatic approach requires a balance in choosing the evolution time: it should be long enough to reduce unwanted diabatic transitions yet short enough to mitigate decoherence from environmental noise. Various approaches have been used to develop quantum shortcut to adiabaticity, where reverse engineering of diabatic transitions is employed to accelerate a slow adiabatic process via nonadiabatic shorcuts [5, 6, 7, 8]. These approaches include counter-diabatic (CD) driving [9, 10, 11, 12, 13], invariant-based approaches [14, 15], the derivative removal of adiabatic gate (DRAG) approach [16], dressed state approaches [17] etc.. Quantum shortcut to adiabaticity has been proposed and experimentally demonstrated in many quantum systems, such as superconducting qubits, defect qubits, quantum dot arrays, and cold atoms [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. One hurdle that prevents the general application of these quantum shortcut approaches is the complexity of the Hamiltonian required for reverse engineering, which often involves multipartite or nonlocal interactions. A number of approaches have been explored to overcome this hurdle by deriving local CD driving, including approach that explores the self-similar nature of a quantum system, mean-field approximation, variational approach, Floquet engineering, digitalized approach, and optimal control approach [31, 32, 33, 34, 38, 35, 36, 37, 39, 40].
Jaynes-Cummings (JC) lattices [41, 42, 43, 44, 45, 46, 47, 48, 49] have been studied in various experimental systems, including superconducting qubits coupled to cavities, solid-state defects coupled to nanocavities, and trapped ions [50, 51, 52, 53, 54, 55, 56, 57], and can exhibit quantum or dissipative phase transitions between the Mott-insulating and superfluid phases in the thermodynamic limit [58, 59, 47, 48]. The states of finite-sized JC lattices can also carry related features, resulting in novel phenomena such as the photon blockade effect and novel entanglement when properly populated with polariton excitations [60, 61]. A key step in demonstrating these effects is to prepare desired quantum states with a finite number of excitations. In previous works, we employed an adiabatic approach and a quantum optimal control approach for state preparation in this system [62, 63].
In our current work, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized JC lattice by applying CD driving during an adiabatic evolution. In our approach, we explore the symmetry of the system’s eigenstates to convert the CD driving to an implementable Hamiltonian for two-site and three-site lattices. We find that for a two-site lattice with one excitation, the CD driving can involve only local qubit-cavity couplings, and with two excitations, the CD driving can have a partial analytical form that involves four-operator couplings in the antisymmetric subspace. We also study the effects of control errors and environmental noise on our scheme, which complements previous CD driving studies on noisy and open quantum systems [36, 64, 65]. Our numerical results demonstrate that the effects of circuit errors and environmental noise are negligible under practical parameters. Meanwhile, this scheme can be characterized by performing measurement on the qubits. This work hence presents an implementable CD driving that can greatly shorten the evolution time of an adiabatic process in a JC lattice and provides a promising avenue to achieving high-fidelity state preparation in a significantly reduced timescale.
II Jaynes-Cummings Lattice
Consider a finite-sized JC lattice composed of JC models coupled via photon hopping between adjacent sites, as illustrated in Fig. 1. The Hamiltonian of the JC lattice can be written as $H_{\rm JC}=H_{0}+gV_{g}+JV_{J}$, where $H_{0}=\sum_{j}(\omega_{0}a_{j}^{\dagger}a_{j}+\frac{\omega_{z}}{2}\sigma_{jz})$ is the Hamiltonian of the uncoupled qubits and cavity modes, $V_{g}=\sum_{j}(a_{j}^{\dagger}\sigma_{j-}+\sigma_{j+}a_{j})$ describes the onsite qubit-cavity coupling, and $V_{J}=-\sum_{j}(a_{i}^{\dagger}a_{i+1}+a_{i+1}^{\dagger}a_{i})$ describes the photon hopping between neighboring sites. Here, $a_{j}$ ($a_{j}^{\dagger}$) is the annihilation (creation) operator of the cavity modes, $\sigma_{j\pm}$, $\sigma_{jz}$ are the Pauli operators of the qubits, $\omega_{0}$ is the frequency of the cavities, $\omega_{z}$ is the qubit energy splitting, $g$ is the strength of the qubit-cavity coupling, $J$ is the photon hopping rate, and $j\in[1,\,N]$ with $N$ being the number of sites. We have assumed $\hbar=1$ for convenience of discussion. In the rotating frame defined by $H_{0}^{\rm rot}=\omega_{z}\sum_{j}(a_{j}^{\dagger}a_{j}+\frac{1}{2}\sigma_{jz})$, the Hamiltonian of the JC lattice becomes $H_{\rm r}=\sum_{j}\Delta a_{j}^{\dagger}a_{j}+gV_{g}+JV_{J}$, where $\Delta=\omega_{0}-\omega_{z}$ is the detuning between the cavity modes and the qubits.
Figure 1: Schematic of a two-site JC lattice with (a) one polariton excitation and (b) two polariton excitations. Each site is composed of a qubit (two-level system) coupled to a cavity mode with coupling strength $g$, and the cavities are connected by photon hopping with rate $J$. In (a), the CD driving includes only local qubit-cavity couplings denoted by $g^{\prime}$; in (b), the CD driving is shown as $H_{1}^{\prime}$.
We denote the basis states of a single JC model as $|n,s\rangle$, with $n\geq 0$ being the photon number of the cavity mode and $s=g,\,e$ being the ground and excited states of the qubit, respectively. The eigenstates of a single JC model include the ground state $|0,g\rangle$ and the doublets $|n,\pm\rangle$ ($n\geq 1$), which are superpositions of the basis states $|n,g\rangle$ and $|n-1,e\rangle$ and correspond to eigenstates with $n$ polariton excitations. In the thermodynamic limit with $N\rightarrow\infty$, a JC lattice can exhibit various many-body effects such as the quantum and dissipative phase transitions between the Mott insulating and the superfluid phases [41, 42, 43]. It can be shown that the ground states of a finite-size JC lattice exhibit similar features to these phases [47, 62, 63].
A key step in studying novel phenomena in a JC lattice is the preparation of desired quantum states for a finite number of polariton excitations. One approach to generate such states is adiabatic evolution [62], where the system is first prepared in the ground state of an initial Hamiltonian $H_{\rm r}(t_{0})$ at time $t_{0}$, followed by adiabatic tuning of the system parameters to reach the target Hamiltonian $H_{\rm r}(T)$ at the final time $T$. The ground state of the target Hamiltonian is designed to be the desired state, which is often hard to prepare directly by gate operations. In the end of the adiabatic evolution, the system will reach the desired state. However, the adiabaticity of the evolution and hence the fidelity of the prepared states can be impaired by unwanted diabatic transitions to excited states. Below, we will employ a quantum shortcut to adiabaticity approach to study the robust generation of quantum states in a JC lattice within a significantly reduced timeframe.
III Counter-Diabatic Driving
CD driving can be applied to the Hamiltonian $H_{\rm r}(t)$ to cancel diabatic transitions and generate quantum shortcut to adiabaticity during an adiabatic evolution. The CD driving has the general form [9, 10, 5]:
$H_{1}(t)=i\sum_{m\neq n}\frac{|m(t)\rangle\langle m(t)|\partial_{t}H_{\rm r}|n(t)\rangle\langle n(t)|}{E_{n}(t)-E_{m}(t)},$ (1)
where $|m\rangle$, $|n\rangle$ are instantaneous eigenstates of $H_{\rm r}(t)$ at time $t$ and $E_{m}$, $E_{n}$ are the corresponding eigenenergies. The total Hamiltonian of this system then becomes $H_{\rm tot}=H_{\rm r}(t)+H_{1}(t)$. The exact CD Hamiltonian in (1) often includes nonlocal or multipartite interactions due to the complexity of the quantum system of interest. Hence, it is challenging to implement this Hamiltonian in systems with practical parameters.
In this section, we convert the CD driving in (1) to an implementable Hamiltonian for adiabatic evolutions in two-site and three-site JC lattices. Our results show that the CD driving can comprise only local qubit-cavity couplings and is implementable using current technology.
III.1 Two sites one excitation
Eigenstates. We first study a two-site JC lattice ($N=2$) with only one polariton excitation. The allowable Hilbert space for this system includes four basis states: $|1,g\rangle_{1}|0,g\rangle_{2}$, $|0,e\rangle_{1}|0,g\rangle_{2}$, $|0,g\rangle_{1}|1,g\rangle_{2}$, and $|0,g\rangle_{1}|0,e\rangle_{2}$, where the subscripts refer to the sites 1 and 2 in the lattice. Written in terms of this basis set, the Hamiltonian has the form:
$H_{\rm r}=\begin{pmatrix}\Delta&g&-J&0\\ g&0&0&0\\ -J&0&\Delta&g\\ 0&0&g&0\end{pmatrix}$ (2)
for given values of $g,\,J,$ and $\Delta$. The eigenvalues $E_{n}$ ($n\in[1,4]$) of this system are
$E_{1,2}=\frac{1}{2}(\Delta_{J}^{-}\mp\chi_{1}^{-}),\quad E_{3,4}=\frac{1}{2}(\Delta_{J}^{+}\mp\chi_{1}^{+})$ (3)
with $\Delta_{J}^{\pm}=\Delta\pm J$ and $\chi_{1}^{\pm}=\sqrt{(\Delta_{J}^{\pm})^{2}+4g^{2}}$. The corresponding eigenvectors are
$v_{1}$ $=\frac{1}{2}\begin{pmatrix}-\sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ -\sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\end{pmatrix},\quad v_{2}=\frac{1}{2}\begin{pmatrix}\sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}+\Delta_{J}^{-}}{\chi_{1}^{-}}}\\ \sqrt{\frac{\chi_{1}^{-}-\Delta_{J}^{-}}{\chi_{1}^{-}}}\end{pmatrix},$ (4)
$v_{3}$ $=\frac{1}{2}\begin{pmatrix}\sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ -\sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ -\sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ \sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\end{pmatrix},\quad v_{4}=\frac{1}{2}\begin{pmatrix}-\sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ -\sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ \sqrt{\frac{\chi_{1}^{+}+\Delta_{J}^{+}}{\chi_{1}^{+}}}\\ \sqrt{\frac{\chi_{1}^{+}-\Delta_{J}^{+}}{\chi_{1}^{+}}}\end{pmatrix}.$
The eigenstate $v_{1}$ is the ground state with energy $E_{1}$. From (4), we observe that the states $v_{1}$ and $v_{2}$ are symmetric with respect to the exchange of the states on the two sites, while the states $v_{3}$ and $v_{4}$ are antisymmetric and acquire a factor of $-1$ when the states on the two sites are swapped.
Figure 2: Eigenenergies $E_{n}$ ($n\in[1,4]$) of a two-site JC lattice with one excitation. (a) $E_{n}$ vs hopping rate $J$ for $g=1$ and $\Delta=1$. (b) $E_{n}$ vs coupling constant $g$ for $J=2$ and $\Delta=1$. All parameters are in dimensionless units.
In Fig. 2(a), we plot the energy spectrum of the system as a function of the photon hopping rate $J$ for $g=1$ and $\Delta=1$. At $J=0$, $\Delta_{J}^{\pm}=\Delta$, resulting in $E_{1}=E_{3}$ ($E_{2}=E_{4}$), which indicates that the states $v_{1}$ and $v_{3}$ ($v_{2}$ and $v_{4}$) are degenerate. In Fig. 2(b), we plot the energy spectrum vs the qubit-cavity coupling $g$ for $J=2$ and $\Delta=1$. At $g=0$ and $\Delta<J$, $E_{2}=E_{3}=0$, and $v_{2}$ and $v_{3}$ are degenerate.
We want to emphasize that the symmetry of these eigenstates is crucial for our development of an implementable CD driving that only involves local qubit-cavity couplings. During an adiabatic evolution, the variation of the Hamiltonian does not induce transition of the system from a symmetric to an antisymmetric state due to the inherent symmetry of the Hamiltonian. Furthermore, this symmetry also facilitates the preparation of the ground state $v_{1}$ at $J=0$ via resonant pumping, without inducing excitation to the state $v_{3}$ that is degenerate to $v_{1}$.
Ramping of hopping rate $J$. Assuming that the hopping rate $J(t)$ is continuously tuned from $J(0)=0$ at time $t=0$ to $J(T)=J_{f}$ at the final time $T$ with $J_{f}$ being the target hopping rate, while keeping the qubit-cavity coupling and the detuning unchanged during the adiabatic process. At $t=0$, the system is prepared in the ground state $v_{1}$ for $J=0$ by resonantly pumping the system from the state $|0,g\rangle_{1}|0,g\rangle_{2}$. The variation of $J(t)$ induces off-diagonal matrix elements between the instantaneous eigenstates, resulting in undesired transitions to the excited states. The nonzero matrix elements are: $|\langle v_{1}|V_{J}|v_{2}\rangle|=g/\chi_{1}^{-}$, $|\langle v_{3}|V_{J}|v_{4}\rangle|=g/\chi_{1}^{+}$, along with their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$, and between the states $v_{3}$ and $v_{4}$, respectively. We then derive the CD driving using (1):
$H_{1}(t)$ $=$ $i\frac{g}{(\chi_{1}^{-})^{2}}\frac{dJ}{dt}(|v_{1}\rangle\langle v_{2}|-|v_{2}\rangle\langle v_{1}|)$ (5)
$-i\frac{g}{(\chi_{1}^{+})^{2}}\frac{dJ}{dt}(|v_{3}\rangle\langle v_{4}|-|v_{4}\rangle\langle v_{3}|).$
Hence, when the system is in the ground state $v_{1}$, the only allowable transition is to the symmetric eigenstate $v_{2}$. As discussed above, this is due to the symmetry of the eigenstates and the Hamiltonian. A variation of the hopping rate in the Hamiltonian preserves the exchange symmetry of the state, which can only induce diabatic transition to the state $v_{2}$. Consequently, the state at an arbitrary time $t$ during the evolution remains symmetric and can only be a superposition of the instantaneous $v_{1}$ and $v_{2}$ states. Thus, the second term in $H_{1}(t)$ does not actively contribute to the elimination of diabatic transitions, and its amplitude can be adjusted with flexibility, without affecting the adiabatic dynamics of this system. Below, we will construct an easier-to-implement CD driving using this property.
We first convert the CD driving in (5) to physical operators of the qubits and cavities to see if it can be easily implemented. In the Hilbert space for one polariton excitation, we have $a_{1}^{\dagger}\sigma_{1-}=|1,g\rangle_{1}\langle 0,e|\otimes|0,g\rangle_{2}\langle 0,g|$, $a_{1}^{\dagger}a_{2}=|1,g\rangle_{1}\langle 0,g|\otimes|0,g\rangle_{2}\langle 1,g|$, and so forth. Using the expression of the eigenstates in (4), we derive:
$|v_{1}\rangle\langle v_{2}|-|v_{2}\rangle\langle v_{1}|=$ $\frac{1}{2}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right),$ (6a)
$|v_{3}\rangle\langle v_{4}|-|v_{4}\rangle\langle v_{3}|=$ $\frac{1}{2}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right),$ (6b)
with $A_{\pm}=a_{1}\pm a_{2}$ and $S_{\pm}=\sigma_{1-}\pm\sigma_{2-}$. And the CD driving is
$H_{1}(t)$ $=$ $i\frac{g}{2(\chi_{1}^{-})^{2}}\frac{dJ}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (7)
$-i\frac{g}{2(\chi_{1}^{+})^{2}}\frac{dJ}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
This Hamiltonian comprises local couplings between qubits and cavities on the same site such as $\sigma_{1+}a_{1}$ and $\sigma_{2+}a_{2}$ as well as nonlocal couplings between qubits and neighboring cavities such as $\sigma_{1+}a_{2}$ and $\sigma_{2+}a_{1}$.
By changing the coefficient of the second term in (7) from $-ig/2(\chi_{1}^{+})^{2}$ to $ig/2(\chi_{1}^{-})^{2}$, we obtain the following CD Hamiltonian:
$H_{1}^{\prime}(t)=ig^{\prime}\left(a_{1}\sigma_{1+}-a_{1}^{\dagger}\sigma_{1-}+a_{2}\sigma_{2+}-a_{2}^{\dagger}\sigma_{2-}\right).$ (8)
This Hamiltonian includes only local qubit-cavity couplings with coupling strength given by $g^{\prime}=\frac{g}{(\chi_{1}^{-})^{2}}\frac{dJ}{dt}$. It can be interpreted as a Jaynes-Cummings coupling with a $e^{i\pi/2}$ phase, and can be implemented by adjusting local qubit-cavity couplings.
For demonstration, we simulate the time evolution of this system under the adiabatic Hamiltonian $H_{\rm r}$ with either $H_{1}$ or $H_{1}^{\prime}$ applied. We consider both linear ramping of the hopping rate with $J(t)=J_{f}\frac{t}{T}$ and quadratic ramping of the hopping rate with $J(t)=J_{f}\left(\frac{t}{T}\right)^{2}$. For linear ramping, $dJ/dt=J_{f}/T$; and for quadratic ramping, $dJ/dt=2J_{f}t/T^{2}$. We choose the target hopping rate $J_{f}=2$, the qubit-cavity coupling $g\equiv 1$, the detuning $\Delta\equiv 1$, and the total time $T=0.5\pi$. We define the fidelity of the quantum state at time $t$ as $F(t)=|\langle v_{1}(t)|v(t)\rangle|^{2}$, where $v(t)$ is the state at time $t$ and $v_{1}(t)$ is the instantaneous ground state of the adiabatic Hamiltonian at time $t$. In Fig. 3(a), we plot the infidelity $1-F(t)$ vs $t/T$. Without the CD driving, the infidelity increases significantly during the evolution, as the diabatic transitions can be serious for a very short evolution time of $T=0.5\pi$. In contrast, when the CD driving $H_{1}^{\prime}$ is applied, the infidelity, up to a small numerical error below $10^{-10}$, is zero throughout the evolution, and the system is preserved in the ground state. In Fig. 3(b), we plot the infidelity $1-F(T)$ of the final state vs the total evolution time $T$. Without the CD driving, the infidelity can be quite high for a short evolution time. This result demonstrates that the CD approach can ensure high fidelity for a short evolution time, which is crucial for devices in the noisy intermediate-scale quantum (NISQ) regime [3]. As shown in Fig. 7(a, b) in Appendix A, the infidelities under $H_{1}$ and $H_{1}^{\prime}$ are the same up to a small numerical error below $10^{-10}$, which confirms our analysis that these two CD drivings yield the same dynamics for the initial state $v_{1}$.
Figure 3: (Color online) (a, c) Infidelity $1-F(t)$ vs $t/T$ for a total evolution time of $T=0.5\pi$ for two-site and three-site JC lattices, respectively. (b, d) Infidelity $1-F(T)$ vs $T$ for two-site and three-site JC lattices, respectively. In all plots, blue solid line (blue circles): linear ramping of hopping rate $J$ for $g\equiv 1$, $\Delta\equiv 1$, $J(0)=0$ and $J_{f}=2$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); red dashed line (red squares): quadratic ramping of $J$ for $g\equiv 1$, $\Delta\equiv 1$, $J(0)=0$ and $J_{f}=2$ under $H_{\rm r}$ (with $H_{1}^{\prime}$); and green dotted line (green triangles): linear ramping of coupling $g$ for $J\equiv 2$, $\Delta\equiv 1$, $g(0)=0$ and $g_{f}=1$ under $H_{\rm r}$ (with $H_{1}^{\prime}$). All parameters are in dimensionless units.
Ramping of qubit-cavity coupling $g$. Now consider an adiabatic trajectory where the qubit-cavity coupling is ramped linearly with $g(t)=g_{f}t/T$ and $g_{f}$ being the target coupling strength. The nonzero transition matrix elements caused by the variation of $g$ are $|\langle v_{1}|V_{g}|v_{2}\rangle|=\Delta_{J}^{-}/\chi_{1}^{-}$, $|\langle v_{3}|V_{g}|v_{4}\rangle|=\Delta_{J}^{+}/\chi_{1}^{+}$, and their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$ as well as between the states $v_{3}$ and $v_{4}$, respectively, due to the symmetry of the eigenstates. The CD driving can be derived as:
$H_{1}(t)$ $=$ $i\frac{\Delta_{J}^{-}}{2(\chi_{1}^{-})^{2}}\frac{dg}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (9)
$-i\frac{\Delta_{J}^{+}}{2(\chi_{1}^{+})^{2}}\frac{dg}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
Similar to (7), this Hamiltonian includes nonlocal couplings between qubits and neighboring cavities. As discussed above, when $|v_{1}\rangle$ is the initial state, only the state $|v_{2}\rangle$ could be excited by diabatic transitions. Therefore, we can vary the coefficient of the second term in (9). By changing this coefficient from $-i\Delta_{J}^{+}/2(\chi_{1}^{+})^{2}$ to $i\Delta_{J}^{-}/2(\chi_{1}^{-})^{2}$, we obtain the CD driving:
$H_{1}^{\prime}(t)=ig^{\prime}\left(a_{1}\sigma_{1}^{+}-a_{1}^{\dagger}\sigma_{1}^{-}+a_{2}\sigma_{2}^{+}-a_{2}^{\dagger}\sigma_{2}^{-}\right),$ (10)
which only involves local qubit-cavity couplings with strength $g^{\prime}=\frac{\Delta_{J}^{-}}{(\chi_{1}^{-})^{2}}\frac{dg}{dt}$. Consequently, implementing both the adiabatic Hamiltonian and the CD driving only requires the tuning of the local qubit-cavity coupling, which can significantly simplify the experimental setup.
We perform a numerical simulation on this system with hopping rate of $J\equiv 2$, detuning of $\Delta\equiv 1$, and a target coupling strength of $g_{f}=1$. At $t=0$ and $g(0)=0$, the initial state is $(|1,g\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,g\rangle_{2})/\sqrt{2}$, where the excitation is stored in the cavity modes. The infidelity $1-F(t)$ vs $t/T$ for $T=0.5\pi$ is plotted in Fig.3(a), and the infidelity at the final time $1-F(T)$ vs the total time $T$ is plotted in Fig.3(b). The results are similar to the case of ramping the hopping rate $J$.
Self-protected trajectory. The magnitude of the CD driving in (10) is proportional to $\Delta_{J}^{-}=\Delta-J$. In the special case of $\Delta=J$, $\Delta_{J}^{-}=0$, leading to the interesting result: $H_{1}^{\prime}=0$, which means that no CD driving is required. It can be shown that the matrix element $|\langle v_{1}|V_{g}|v_{2}\rangle|\equiv 0$ under this condition, hence the only allowable diabatic transition for the initial state $v_{1}$ disappears. The Hamiltonian $H_{\rm r}$ is thus self-protected from diabatic transition, i.e., there is no diabatic transition when ramping the qubit-cavity coupling $g$. The ground state is $\frac{1}{\sqrt{2}}\left(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2}\right)$ throughout the evolution, with the system occupying the lower polariton state $|1,-\rangle$ in one of the lattice sites.
III.2 Three sites one excitation
[…]
$H_{\rm r}=\begin{pmatrix}\Delta&g&-J&0&0&0\\ g&0&0&0&0&0\\ -J&0&\Delta&g&-J&0\\ 0&0&g&0&0&0\\ 0&0&-J&0&\Delta&g\\ 0&0&0&0&g&0\end{pmatrix}.$ (11)
[…]
$H_{1}$ $=$ $ig_{n}\left[\left(a_{1}^{\dagger}+a_{3}^{\dagger}\right)\sigma_{2-}+a_{2}^{\dagger}\left(\sigma_{1-}+\sigma_{3-}\right)-h.c.\right]$ (12)
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}-h.c.\right)$
$+ig_{l}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}-h.c.\right)$
which includes nonlocal couplings with strengths $g_{n}$ and $g_{l}$, such as the coupling between the cavities on sites $1$ and $3$ and the qubit on site $2$, as well as local couplings with strength $g_{l}$ for sites 1, 3131,\,31 , 3 and strength $2g_{l}$ for site $2$.
We employ the same approach as in Sec. III.1 to convert the CD Hamiltonian to a form that is easier to implement. As discussed above, for initial state $v_{1}$, ramping of the hopping terms only induces transition to the state $v_{4}$. Therefore, the coefficient $\alpha_{2}$ in $H_{1}$ can be varied with flexibility without affecting the dynamics of this system. By adjusting $\alpha_{2}$, we find
$H_{1}^{\prime}$ $=$ $ig^{\prime}\left(a_{1}^{\dagger}\sigma_{1-}+2a_{2}^{\dagger}\sigma_{2-}+a_{3}^{\dagger}\sigma_{3-}-h.c.\right)$ (13)
$+ig^{\prime}\left(a_{1}^{\dagger}\sigma_{3-}+a_{3}^{\dagger}\sigma_{1-}-h.c.\right),$
where the couplings between the qubit (cavity) on site $2$ and the cavities (qubits) on sites $1$ and $3$ are no longer required. This Hamiltonian still includes nonlocal terms between sites $1$ and $3$ with coupling strength $g^{\prime}$. This Hamiltonian can be implemented for a small lattice of three sites by coupling the qubits and cavities on sites $1$ and $3$. However, this result shows that we cannot always find a local form of the CD driving. In Fig. 3(c, d), we plot the infidelity of the prepared state under the adiabatic Hamiltonian $H_{\rm r}$ both without and with the CD driving $H_{1}^{\prime}$. The result is similar to that of the two-site lattice discussed in Sec. III.1.
III.3 Two sites two excitations
Eigenstates. For a two-site JC lattice with two polariton excitations, the Hilbert space includes eight basis states: $|2,g\rangle_{1}|0,g\rangle_{2}$, $|1,e\rangle_{1}|0,g\rangle_{2}$, $|0,g\rangle_{1}|2,g\rangle_{2}$, $|0,g\rangle_{1}|1,e\rangle_{2}$, $|1,g\rangle_{1}|1,g\rangle_{2}$, $|0,e\rangle_{1}|1,g\rangle_{2}$, $|1,g\rangle_{1}|0,e\rangle_{2}$, and $|0,e\rangle_{1}|0,e\rangle_{2}$. In the first four states, both excitations occupy one of the two sites, while in the last four states, one excitation is located on each site. Using this basis set, the Hamiltonian $H_{\rm r}$ can be expressed as:
$H_{\rm r}=\begin{pmatrix}2\Delta&\sqrt{2}g&0&0&-\sqrt{2}J&0&0&0\\ \sqrt{2}g&\Delta&0&0&0&-J&0&0\\ 0&0&2\Delta&\sqrt{2}g&-\sqrt{2}J&0&0&0\\ 0&0&\sqrt{2}g&\Delta&0&0&-J&0\\ -\sqrt{2}J&0&-\sqrt{2}J&0&2\Delta&g&g&0\\ 0&-J&0&0&g&\Delta&0&g\\ 0&0&0&-J&g&0&\Delta&g\\ 0&0&0&0&0&g&g&0\end{pmatrix}.$ (14)
For detuning $\Delta=0$, the eigenvalues of this Hamiltonian can be derived analytically as:
$E_{1,2}$ $=0,$ (15a)
$E_{3,4}$ $=\mp\sqrt{2g^{2}+J^{2}},$ (15b)
$E_{5,6}$ $=\mp\frac{\sqrt{6g^{2}+5J^{2}-\sqrt{4g^{4}+60g^{2}J^{2}+9J^{4}}}}{\sqrt{2}},$ (15c)
$E_{7,8}$ $=\mp\frac{\sqrt{6g^{2}+5J^{2}+\sqrt{4g^{4}+60g^{2}J^{2}+9J^{4}}}}{\sqrt{2}}.$ (15d)
In Fig. 4(a), we plot the eigenenergies vs the hopping rate $J$ at $g=1$ and $\Delta=0$. We denote the eigenstates as $v_{i}$ with $i\in[1,8]$, corresponding to the eigenvalues given in (15). Here, $v_{7}$ is the ground state, while $v_{1}$ and $v_{2}$ are degenerate states with $E_{1,2}=0$.
Figure 4: (Color online) (a) Eigenenergies $E_{n}$ ($n\in[1,8]$) vs hopping rate $J$. (b) Infidelity $1-F(t)$ vs $t/T$ for a total evolution time of $T=0.5\pi$. Blue solid line (blue circles): linear ramping of hopping rate $J$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}$ applied) with the initial state being the ground state $v_{7}$ at $J=0$; and red dashed line (red squares): linear ramping of $J$ under $H_{\rm r}$ (with $H_{1}^{\prime}$) with the initial state being the first excited state $v_{3}$ at $J=0$. Here, $g\equiv 1$, $\Delta\equiv 0$, $J(0)=0$, $J_{f}=1$, and all parameters are in dimensionless units.
Ramping of hopping rate $J$. Assuming that the hopping rate is linearly ramped with $J=J_{f}t/T$, the target hopping rate $J_{f}=1$, $g\equiv 1$, and $\Delta\equiv 0$. The CD driving $H_{1}$ can be obtained numerically using Eq. (1). We simulate the dynamics of this system with the initial state being the ground state $v_{7}$ at $J=0$ for a total evolution time of $T=0.5\pi$. In Fig. 4(b), we plot the infidelity $1-F(t)$ of the quantum state at time $t$ as a function of $t/T$. The infidelity at the final time $T$ exceeds $0.4$ under the adiabatic Hamiltonian $H_{\rm r}$, whereas it remains negligible when the CD driving is applied.
By examining the eigenstates of $H_{\rm r}$, we find that the eight eigenstates can be divided into two subsets. The first subset contains three states $\{v_{2},\,v_{3},\,v_{4}\}$; while the second subset contains the remaining five states, including the ground state $v_{7}$. When varying the hopping rate $J$, the only nonzero transition matrix elements are between states within the same subset, and there is no diabatic transition between states in different subsets. Thus, for an initial state in the first subset, we consider the CD driving $H_{1}^{\prime}$ that only eliminates the diabatic transitions between eigenstates in this subset. As detailed in Appendix B, we derive
$H_{1}^{\prime}(t)=i\frac{g}{2\left(J^{2}+2g^{2}\right)}\frac{dJ}{dt}\left(A_{2}^{\dagger}C_{2}-C_{2}^{\dagger}A_{2}\right),$ (16)
where $A_{2}=a_{1}^{2}-a_{2}^{2}$ and $C_{2}=a_{2}\sigma_{1-}-a_{1}\sigma_{2-}$. This Hamiltonian consists of four-operator terms, which can be formed through effective second-order couplings in the perturbative regime.
To demonstrate the effectiveness of Eq. (16), we conducted a numerical simulation of the adiabatic process on the initial state $v_{3}$ from the first subset, which corresponds to the first excited state in Fig. 4(a). As shown in Fig. 4(b), up to a small numerical error below $10^{-10}$, the infidelity remains zero throughout the evolution, indicating that the diabatic transitions are completely cancelled by the CD driving $H_{1}^{\prime}$. In contrast, if the initial state is the ground state $v_{7}$ from the second subset, the infidelities both without and with the CD driving $H_{1}^{\prime}$ will be the same [see the blue solid line in Fig. 4(b)]. These numerical findings clearly show that the CD driving $H_{1}^{\prime}$ is only effective for states in the first subset, but not for states in the second subset, which is consistent with our analysis above.
IV Error, Noise, and Measurement
In this section, we discuss the effects of control errors and environmental noise on the adiabatic evolution under the CD driving, as well as the characterization of the CD approach through qubit measurements.
Control errors. Adjusting the system parameters can introduce classical control errors, potentially impacting the performance of the CD approach. To investigate the effects of these errors, we conducted numerical simulations of the adiabatic process by introducing random fluctuations to the time-dependent coupling constants. Consider the linear ramping of the hopping rate $J(t)$ discussed in Sec. III.1, where the CD driving only consists of time-dependent local qubit-cavity couplings. We introduce a fluctuation $\delta J(t)$ to the hopping rate with $J(t)\rightarrow J(t)+\delta J(t)$, and a fluctuation $\delta g(t)$ to the qubit-cavity coupling [including the CD terms in (10)] with $g(t)\rightarrow g(t)+\delta g(t)$. Here, $\delta J(t)$ and $\delta g(t)$ are Gaussian random numbers with a standard deviation $\alpha J_{f}$ and $\alpha g$, respectively, where $\alpha$ is the ratio of the fluctuation amplitude to the coupling strength. We choose $\alpha$ to be below $0.15$, i.e., the fluctuations are below $15\%$ of the coupling strengths. In superconducting quantum devices, the couplings can be controlled by various methods, as demonstrated in recent experiments [66, 67, 68, 69, 70, 71, 72], and the ratio $\alpha$ can be well below $5\%$ [73, 74]. In Fig. 5(a), the fidelity $F(t)$ is plotted vs $t/T$ for a single sample of errors with $\alpha=0.05$ and $T=0.5\pi$. The fidelity exhibits small fluctuations over the time $t$ due to the presence of the errors. In Fig. 5(b), the fidelity $F(T)$ averaged over $100$ samples of errors is plotted vs the ratio $\alpha$. As $\alpha$ increases to $15\%$, the fidelity only shows a mild decrease, demonstrating that control errors will not strongly affect the system dynamics in practical systems.
Figure 5: (Color online) (a) Fidelity $F(t)$ vs $t/T$ for one sample of errors with $\alpha=0.05$. (b) Fidelity $F(T)$ vs $\alpha$ averaged over 100 samples of errors. In (a) and (b), blue circles (red triangles): under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); and the total evolution time is $T=0.5\pi$. (c) Infidelity $1-F(T)$ vs $T$ for $\gamma=\frac{5}{\pi}\times 10^{-5}$. Blue solid line (blue circles): for $\kappa=5\times 10^{-4}$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); red dashed line (red squares): for $\kappa=10^{-4}$ under $H_{\rm r}$ (with $H_{1}^{\prime}$); and green dotted line (green triangles): for $\kappa=5\times 10^{-5}$ under $H_{\rm r}$ (with $H_{1}^{\prime}$). (d) Infidelity $1-F(T)$ vs $\kappa$ for $\gamma=\frac{5}{\pi}\times 10^{-5}$. Blue solid line (blue circles): for $T=0.5\pi$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); red dashed line (red squares): for $T=3\pi$ under $H_{\rm r}$ (with $H_{1}^{\prime}$); and green dotted line (green triangles): for $T=5.5\pi$ under $H_{\rm r}$ (with $H_{1}^{\prime}$). All plots are for linear ramping of hopping rate $J$ with parameters given in Fig. 3 in dimensionless units.
Decoherence. Environmental noise induces decoherence in the qubits and the cavity modes, potentially impacting the fidelity of the prepared quantum states. To analyze the effects of decoherence, we utilize a master equation approach with
$\frac{d\rho}{dt}=-i\left[H_{\rm tot},\rho\right]+\sum_{j}\left(\gamma_{j}\mathcal{L}_{qj}+\kappa_{j}\mathcal{L}_{aj}\right)\rho,$ (17)
where $\mathcal{L}_{qj}=\frac{1}{2}(2\sigma_{j-}\rho\sigma_{j+}-\rho\sigma_{j+}\sigma_{j-}-\sigma_{j+}\sigma_{j-}\rho)$ is the Liouvillian operator for the qubit on site $j$ with damping rate $\gamma_{j}$, $\mathcal{L}_{aj}=\frac{1}{2}(2a_{j}\rho a_{j}^{\dagger}-\rho a_{j}^{\dagger}a_{j}-a_{j}^{\dagger}a_{j}\rho)$ is the Liouvillian operator for the cavity mode on site $j$ with damping rate $\kappa_{j}$, and $H_{\rm tot}$ is the total Hamiltonian of the system. We choose $\gamma_{1,2}=\gamma$ and $\kappa_{1,2}=\kappa$ for simplicity of discussion. We also assume that the dimensionless coupling strength $g=1$ in our discussion corresponds to $g=2\pi\times 100$ MHz in superconducting systems [62, 63]. The measured decoherence time of $\sim 100{\rm\mu s}$ then corresponds to $\gamma=\frac{5}{\pi}\times 10^{-5}$ in dimensionless units. When the cavity frequency is $\omega_{0}=2\pi\times 5$ GHz and the quality factor is $Q=10^{6}$, the cavity decay rate is $\kappa=5\times 10^{-5}$ in dimensionless units.
[…]
Figure 6: (Color online) Operator average $\langle\sigma_{1x}\sigma_{2x}\rangle$ vs $t/T$ for $T=0.5\pi$. (a) Two sites one excitation with linear ramping of $J$ and parameters given in Fig. 3 in dimensionless units. (b) Two sites two excitations with linear ramping of $J$ and parameters given in Fig. 4 in dimensionless units. In (a) and (b), blue solid line (blue circles): under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); and red dashed line: instantaneous ground state average of $H_{\rm r}$.
Measurement. The performance of the CD approach can be characterized by conducting measurement on the qubit operator $\sigma_{1x}\sigma_{2x}$. In the case of one excitation in a two-site lattice, the ground state at the initial parameters $g=1$, $J=0$, and $\Delta=1$ is $\frac{1}{2}(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2})$, where $|1,-\rangle$ is the lower-polariton state with one excitation, and the excitation occupies the sites in the form of a quantum superposition with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0.7236$. When the hopping rate becomes $J_{f}=2$ at time $T$, the operator average of the ground state decreases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.2764$, as shown in Fig. 6(a). It is noticeable that the operator average for the adiabatic evolution under $H_{\rm r}$ can significantly deviate from that of the ground state, while the operator average with the CD driving applied remains the same as that of the ground state. Similarly, in the two-excitation case, the ground state at $g=1$, $J=0$, and $\Delta=0$ is $v_{7}=|1,-\rangle_{1}|1,-\rangle_{2}$, where each site is occupied by one excitation in the lower-polariton state with the operator average $\langle\sigma_{1x}\sigma_{2x}\rangle=0$. When the hopping rate becomes $J_{f}=1$, the operator average increases to $\langle\sigma_{1x}\sigma_{2x}\rangle=0.3659$. During the evolution, the operator average for the adiabatic process without the CD driving shows a large discrepancy with the operator average in the CD approach, as shown in Fig. 6(b). Hence, the effectiveness of the CD approach can be verified by measuring the operator average.
V Conclusions
To conclude, we have studied a quantum shortcut to adiabaticity using the counter-diabatic driving approach in a finite-sized JC lattice, leveraging the symmetry inherent in its eigenstates. Our analysis reveals that for a two-site lattice with one excitation, the CD driving can be simplified to a realizable form that only consists of local qubit-cavity couplings. For more complex scenarios, such as a two-site lattice with two excitations or a three-site lattice with one excitation, a simpler CD Hamiltonian can still be obtained, albeit involving nonlocal terms. Our numerical simulations demonstrate that control errors and environmental noise have negligible effects on our scheme in practical systems and our scheme can be characterized through measurements on the qubits. Overall, our findings offer a promising avenue for achieving high-fidelity quantum state preparation and hold potential for applications across various systems exhibiting similar symmetries.
Acknowledgements
This work is supported by the NSF Award No. 2037987 and the UC-MRPI Program (Grant ID M23PL5936). A.G. is also supported by the Sandbox AQ Fellowship.
Appendix A Relation between $H_{1}$ and $H_{1}^{\prime}$
In Sec. III.1, we give an analysis that the system dynamics under the CD driving $H_{1}^{\prime}$ will be exactly the same as that of the CD driving $H_{1}$, when the initial state is $v_{1}$. In Fig. 3(a, b), we plot the infidelity for a two-site lattice with one excitation under the CD driving $H_{1}^{\prime}$, where the infidelity remains zero up to a small numerical error below $10^{-10}$. In Fig. 7(a, b), we plot the infidelity of this system under both $H_{1}$ and $H_{1}^{\prime}$ for a linear ramping of the hopping rate $J$ using the same parameters as given in Fig. 3(a, b). It can be clearly seen that the infidelities for both $H_{1}$ and $H_{1}^{\prime}$ remain zero up to the numerical error, which confirms our analysis.
Figure 7: (Color online) (a) Infidelity $1-F(t)$ vs $t/T$ for total evolution time $T=0.5\pi$. (b) Infidelity $1-F(T)$ vs $T$. In (a) and (b), red solid line (blue circles): under the adiabatic Hamiltonian $H_{\rm r}$ with the CD driving $H_{1}$ ($H_{1}^{\prime}$) applied. (c) Average energy $E_{\rm ave}$ vs $t/T$. (d) Energy cost $\Delta E$ vs $t/T$. In (c) and (d), blue solid line: under $H_{\rm r}$ only; red circles: with $H_{1}$ applied; and blue dots: with $H_{1}^{\prime}$ applied. All the plots are for two sites one excitation with linear ramping of $J$, and the parameters are given in Fig. 3 in dimensionless unit.
Furthermore, we compare the effect of the CD drivings $H_{1}$ and $H_{1}^{\prime}$ by numerically calculating two quantities: (1) the average energy of the system $E_{\rm ave}=\langle v(t)|H_{\rm tot}(t)|v(t)\rangle$ at time $t$ with $H_{\rm tot}$ being the total Hamiltonian and $v(t)$ being the instantaneous wavefunction, and (2) the energy cost $\Delta E=E_{\rm ave}-E_{G}$ with $E_{G}$ being the instantaneous ground state energy at time $t$ following the definition in [75]. In Fig. 7(c), we plot $E_{\rm ave}$ for linear ramping of the hopping rate $J$ on the initial state $|v_{1}\rangle$ under the adiabatic Hamiltonian only, with the CD driving $H_{1}$ applied, and with $H_{1}^{\prime}$ applied, respectively. The average energy for the adiabatic Hamiltonian $H_{\rm tot}=H_{\rm r}$ is higher than that with the CD driving $H_{1}$ or $H_{1}^{\prime}$, indicating nonzero diabatic transition to the excited states, while the average energy with the CD driving $H_{1}$ is the same as that with $H_{1}^{\prime}$, confirming that these two CD Hamiltonians give the same dynamics. In Fig. 7(d), we plot the energy cost $\Delta E$ vs the time $t/T$. For the adiabatic process under the Hamiltonian $H_{\rm tot}=H_{\rm r}$, $\Delta E=0$ at time $t=0$, because the initial state is the ground state of the instantaneous Hamiltonian. During the evolution, $\Delta E$ continuously increases due to diabatic transition to the excited state. Under the CD driving $H_{1}$ or $H_{1}^{\prime}$, the energy cost is nonzero. This is because the state $v(t)$ at an arbitrary time $t$ is the instantaneous ground state of $H_{\rm r}$ but not the ground state of the total Hamiltonian with the CD driving $H_{1}$ or $H_{1}^{\prime}$ being nonzero at any time $t$ including the initial and the final times.
Appendix B CD Driving for Two Excitations
For a two-site lattice with two polariton excitations, the adiabatic Hamiltonian is given in Eq. (14). Solving this Hamiltonian for $\Delta=0$, we can find its eigenvalues and eigenvectors. The eigenvalues are given in (15), and the eigenvectors are
$|v_{1}\rangle$ $=\left(0,-JA_{1},0,-JA_{1},-gA_{1},0,0,1\right),$
$|v_{2}\rangle$ $=\left(A_{2},0,-A_{2},0,0,-1,1,0\right),$
$|v_{3,4}\rangle$ $=\left(-\frac{1}{A_{2}},\pm\frac{1}{A_{3}},\frac{1}{A_{2}},\mp\frac{1}{A_{3}},0,-1,1,0\right),$
$|v_{5,6}\rangle$ =(±$=\left(\pm X_{-},Y_{-},\pm X_{-},Y_{-},Z_{-},\pm M_{-},\pm M_{-},1\right),$
$|v_{7,8}\rangle$ $=\left(\mp X_{+},Y_{+},\mp X_{+},Y_{+},Z_{+},\mp M_{+},\mp M_{+},1\right),$
where we define
$A_{1}$ $=\frac{g}{g^{2}-J^{2}},\,\,A_{2}=-\frac{J}{\sqrt{2}g},\,\,A_{3}=-\frac{J}{\sqrt{2g^{2}+J^{2}}},\,\,A_{4}=\sqrt{4g^{4}+60g^{2}J^{2}+9J^{4}},$
$X_{\pm}$ $=-\frac{24J}{g^{2}}\frac{[\pm 2g^{6}+g^{4}(\pm 19J^{2}+A_{4})+J^{4}(\pm 3J^{2}+A_{4})+g^{2}(\pm 19J^{4}+3J^{2}A_{4})]}{\sqrt{6g^{2}+5J^{2}\pm A_{4}}(\pm 2g^{2}\pm 9J^{2}+A_{4})(\pm 2g^{2}\pm 3J^{2}+A_{4})},$
$Y_{\pm}$ $=-\frac{8J}{g}\frac{[16g^{6}+g^{4}(166J^{2}\pm 8A_{4})\pm 12J^{4}(\pm 3J^{2}+A_{4})+g^{2}(201J^{4}\pm 29J^{2}A_{4})]}{(2g^{2}+3J^{2}\pm A_{4})(6g^{2}+5J^{2}\pm A_{4})(2g^{2}+9J^{2}\pm A_{4})},$
$Z_{\pm}$ $=\frac{2}{g^{2}}\frac{[8g^{6}+g^{4}(90J^{2}\pm 4A_{4})\pm 8J^{4}(3J^{2}+A_{4})+g^{2}(125J^{4}\pm 17J^{2}A_{4})]}{(\pm 2g^{2}\pm 9J^{2}+A_{4})(\pm 6g^{2}\pm 5J^{2}+A_{4})},$
$M_{\pm}$ $=\frac{2\sqrt{2}}{g}\frac{[\pm 2g^{4}+g^{2}(\pm 11J^{2}+A_{4})+J^{2}(\pm 3J^{2}+A_{4})]}{(2g^{2}+3J^{2}\pm A_{4})\sqrt{6g^{2}+5J^{2}\pm A_{4}}}.$
Note that the eigenvectors given above have not been normalized. We find that the eigenvectors can be grouped into two subsets. The first subset is composed of the antisymmetric eigenvectors $\{v_{2},v_{3},v_{4}\}$, and the second subset is composed of the symmetric eigenvectors $\{v_{1},v_{5},v_{6},v_{7},v_{8}\}$.
Using the expressions of the eigenvalues, eigenvectors, and Eq. (1), we can derive the CD driving $H_{1}$ for this system, but we cannot write $H_{1}$ into an analytical form. Neither can we find an implementable CD driving as we did in Sec. III.1 for the case of one excitation. As varying the adiabatic Hamiltonian $H_{\rm r}$ only induces diabatic transitions between eigenstates of the same symmetry, the CD driving only contains nonzero matrix elements between eigenstates in the same subset. We find that when considering only the matrix elements for the first subset $\{v_{2},v_{3},v_{4}\}$, the CD Hamiltonian can be written in terms of the qubit and cavity operators as
$H_{1}^{\prime}=i\frac{g}{2(J^{2}+2g^{2})}\frac{\partial J}{\partial t}\left[(a_{1}^{\dagger})^{2}\sigma_{1-}a_{2}-(a_{1}^{\dagger})^{2}a_{1}\sigma_{2-}-(a_{2}^{\dagger})^{2}\sigma_{1-}a_{2}+(a_{2}^{\dagger})^{2}a_{1}\sigma_{2-}\right]+h.c.$ (20)
Defining $A_{2}=a_{1}^{2}-a_{2}^{2}$ and $C_{2}=a_{2}\sigma_{1-}-a_{1}\sigma_{2-}$, it can be shown that Eq. (20) leads to Eq. (16) in Sec. III.3. On the other hand, we cannot find an analytical expression for the CD driving for the second subset that contains the ground state $v_{7}$.